# Reducing Task Discrepancy of Text Encoders for Zero-Shot Composed Image Retrieval

### v1

@@ 3.3 Can RTD really reduce the task discrepancy of the text encoder?
Moreover, we additionally measure the average cosine similarity between the composed textual features with the prompt “a photo of $\phi(\psi_{V}(I_{r}))$ that $T_{c}$” (Figure 1a) and the target image features (Figure 1c). The similarity is measured by the LinCIR ViT-L/14 model [17] on the CIRCO validation split. When we use the frozen CLIP text encoder ($\psi_{T}$), the average similarity is 0.1. By changing the text encoder to our updated text encoder ($\psi_{T}^{tr}$), the similarity becomes 0.29 (+0.19). This result shows that RTD successfully aligns the composed query features using $\phi$ to the frozen CLIP image features.
4 Experiments
4.1 Experimental setup
Implementation details. We use the AdamW optimizer [28] with a weight decay of 0.01. The learning rate is set to $10^{-5}$, with a batch size of 512. For a fair comparison, we select the text encoder model with the best zero-shot CIRR [1] dev R@1 score for evaluating RTD. We evaluate the CIR performances of the model in a zero-shot manner by evaluating it across four different benchmarks. We use the visual and textual encoders of the CLIP ViT-B/32 and ViT-L/14 [14] as our backbone. Unless otherwise noted, we use the LLM-based 2.5M text triplets provided by CompoDiff [8] for the training. We set the $\tau$ as $0.07$ in Eq. 1 and scale the standard deviation of Gaussian distribution as $0.5$ for the noise injection. More results on various noise distributions can be found in the Appendix. All experiments were conducted using four NVIDIA A100 GPUs with Python 3.8 and Pytorch [29].
Evaluation datasets and metrics. We compare ZS-CIR methods on five benchmark datasets: CIRR [1], CIRCO [16], FashionIQ [18], COCO [30], and GeneCIS [19]. Details of each dataset are in the Appendix A.1. For CIRR, FashionIQ, COCO, and GeneCIS, we have reported their recall scores at the top K retrieval results (R@K). Since the CIRCO dataset includes multiple positive images for each query, we use a ranking-based metric—mean Average Precision scores at the top K results (mAP@K)—which provides a more robust and reliable assessment [31, 32]. For the main results, we compare the results on the three categories (Shirt, Dress, Toptee) of the FashionIQ validation split, as well as the test sets of CIRR and CIRCO. GeneCIS and COCO object composition results and their detailed explanations can be found in the Appendix B.1.
Baselines. We evaluate the effect of our method when combined with three recent ZS-CIR methods: Pic2Word [15], SEARLE [16], and LinCIR [17]. All three methods share the same core concept shown in Figure 1, but use different training schemes. Pic2Word[15] optimizes contrastive loss between the image embedding and its projected text embedding of “a photo of [\$]” to obtain the projection module $\phi$. Similarly, SEARLE [16] employs a two-stage approach, starting with an optimization-based textual inversion phase followed by a distillation phase for the projection module $\phi$. LinCIR [17] introduces a language-only self-supervised task involving keyword token replacement by letting the original text embedding and the replaced text embedding whose keyword tokens are changed to the projected original text embedding by $\phi$.
We train all these methods with the same backbone architectures (CLIP ViT-B/32 and ViT-L/14). We use the publicly available pre-trained model for SEARLE (ViT-B/32, ViT-L/14) and Pic2Word (ViT-L/14). Otherwise, we reproduce the results using the official implementation. When reproducing, we adhere to the same settings in the original papers. For example, we select the final last epoch model for the Pic2Word ViT-B/32 model and choose the model based on the best zero-shot CIRR [1] dev R@1 score for LinCIR. Although our method is not specifically designed for projection-based method, we do not compare our method with CIR methods that require massive external triplet datasets [8, 9, 11] or those utilizing large-scale captioners and LLMs during inference [13, 33] due to their inefficiency. Integrating RTD into the other CIR methods will be an interesting future work.

### v2

@@ 4.4 Anaylses on our core motivation
We additionally measure the average cosine similarity between the composed textual features with the prompt “a photo of $\phi(\psi_{V}(I_{r}))$ that $T_{c}$” (Fig. 1a) and the target image features (Fig. 1c). The similarity is measured by the LinCIR ViT-L/14 model on the CIRCO validation split. When we use the frozen CLIP text encoder ($\psi_{T}$), the average similarity is 0.1. By changing the text encoder to our updated text encoder ($\psi_{T}^{tr}$), the similarity becomes 0.29 (+0.19). This result shows again that RTD successfully aligns the composed query features using $\phi$ to the frozen CLIP image features.
Is all the gain of RTD from naive text-encoder updating? To verify that our improvements cannot be achieved solely by updating the text encoder backbone without considering the task discrepancy, we additionally measure the results of previous methods (Pic2Word and LinCIR) when naively updating the text encoders. Namely, after training $\phi$ while keeping all other networks frozen as in previous methods, we additionally update the text encoder using the original loss function, while fixing other modules including $\phi$. We denote this update rule as “naïve” in the Tab. 5. Unlike RTD, we observe that just naively updating the text encoder (“naïve”) significantly degrades the performance of the baseline. The results indicate that merely updating the text backbone is not beneficial for CIR; instead, mitigating task discrepancy through RTD is necessary.
Table 5: Impact of the update scheme. Two update schemes are compared: (1) using the original objective from baseline (“naïve”) and (2) using RTD. For a fair comparison, in both schemes, $\phi$ is updated first and $\psi_{T}$ is updated top on the frozen modules. Other details are the same as Tab. 3.
CIRR (R) CIRCO (mAP) FIQ (R) Avg
@5 @10 @10 @25 @10 @50
Pic2Word 51.4 64.4 8.8 10.1 25.3 44.9 32.2
+naïve 19.2 27.5 1.3 1.6 4.4 11.2 10.9
+ RTD 56.6 69.8 8.8 9.8 27.6 48.9 36.9
LinCIR 54.3 67.8 12.7 14.5 27.4 47.7 36.9
+naïve 52.7 66.8 11.4 13.0 26.3 45.9 35.5
+ RTD 57.9 71.1 16.1 17.8 30.2 51.1 40.7
Next, instead of using the original objective from baselines, we update the text encoder with conventional contrastive learning using the pair $(T_{r},T_{r})$, which corresponds exactly to row 2 in Tab. 3 and row 3 in Tab. 4. This setup can also serve as a proxy for the standalone impact of breaking the alignment of pre-trained knowledge, as noted in [3]. As demonstrated in Sec. 4.3, the large gap (more than 2 points) between EPC (last row in Tab. 3) and this proxy (row 2 in Tab. 3) again shows that our gain cannot be attributed to this simple alternative (naive text encoder tuning).
Figure 3: Impact of sizes of backbone. The results of RTD combined with Pic2Word [36], SEARLE [4], and LinCIR [17] across different CLIP backbones (ViT-B/32 and ViT-L/14) are shown. Here, the score is the same metric in “Avg” in Tab. 3 and other details are the same as Tab. 3. Full results are in the Sec. B.2.
4.5 Compatibility across backbone sizes
In the Fig. 3, we observe that the integration of our approach with projection-based CIR methods significantly improves the performance across pioneering projection-based ZS-CIR methods (SEARLE, Pic2Word, and LinCIR) and all backbones (ViT-B/32 and ViT-L/14). For example, regardless of the choice of projection module $\phi$ and backbones, the minimum performance gain for average scores is greater than 2.8 points. The performance of RTD using the larger backbone (ViT-G/14) can be found in Tab. 2. Details and the full results are provided in the Sec. B.3. We also provide additional qualitative retrieval results in the App. D.
Table 6: The effectiveness of different types of text triplets for RTD. “In-context” denotes an efficient implementation using in-context learning with LLaMA3-8B, without fine-tuning. Details and examples of each text triplet dataset are summarized in Tab. 7 and Tab. 8, respectively. Other details are the same as Tab. 3.
Source LLM CIRR (R) CIRCO (mAP) FIQ (R) Avg
@5 @10 @10 @25 @10 @50
LinCIR - - 54.3 67.8 12.7 14.5 27.4 47.7 37.4
+RTD Rule-based 56.7 70.3 15.0 17.0 30.4 51.9 40.2 (+ 2.9)
IP2P [5] 58.7 71.6 15.9 18.0 29.6 50.7 40.7 (+ 3.3)
Compodiff [16] 57.9 71.1 16.1 17.8 30.2 51.1 40.7 (+ 3.3)
In-context 59.3 71.8 15.8 17.5 29.7 51.4 40.9 (+ 3.5)
CoVR [45] 59.8 72.6 15.4 17.0 29.6 50.8 41.9 (+ 4.5)
CASE [22] 56.3 69.3 11.1 12.7 26.6 47.8 37.7 (+ 0.3)
4.6 Impact of the text triplet generation strategies
As explained in Sec. 3.1, we evaluate RTD using both 1) publicly available LLM-based text triplets (from IP2P, Compodiff, CoVR, and CASE) along with efficient in-context learning-based text triplets, and 2) LLM-free, rule-based triplets. Tab. 6 shows that RTD consistently improves CIR performance across them (+3.3 for IP2P, +3.3 for Compodiff, +3.5 for in-context learning, +4.5 for CoVR, and +0.31 for CASE, and rule-based triplets achieve 2.9, respectively). We believe this result shows the reproducibility and consistency of RTD, with the rule-based triplets performing comparably to LLM-generated ones, indicating that efficient rule-based triplets are sufficient to achieve strong CIR performance. The marginal improvement in CASE is largely due to the poor quality of text triplets resulting from its construction pipeline that prioritizes CIR triplet quality over text triplet quality, as shown in Tab. 7. Further details can be found in Sec. A.2, and additional analyses, including data scales related to text triplets, are provided in Sec. B.5.
5 Conclusion
We presented RTD, a novel post-processing approach that can be easily integrated into existing projection-based CIR methods, aimed at reducing task discrepancy of text encoders. Empirical evaluations demonstrate that RTD significantly boosts the performance of existing projection-based CIR methods across diverse datasets and model backbones, competing with or outperforming other resource-intensive CIR methods with much greater efficiency.
Supplementary Material
Appendix A Additional Implementation Details
A.1 CIR Datasets
We additionally provide results on two more benchmark datasets, GeneCIS [44] and COCO Object Composition introduced by [36], in Sec. B.1. GeneCIS [44] is also constructed based on COCO images and the Visual Attributes in the Wild dataset [33]. GeneCIS introduces four task variations: (1) focus on an attribute, (2) change an attribute (3) focus on an object, and (4) change an object. These tasks explore different aspects of image retrieval and manipulation. For the COCO Object Composition task, we utilize 5000 images from the COCO validation dataset to evaluate object composition. Our objective is to retrieve an image that contains an object specified by a query image, along with scenes or objects described using text. The composed query is constructed by combining “a photo of [\$], [$obj_{1}$], [$obj_{2}$] … and [$obj_{n}$]” where [$obj_{i}$] are text descriptions.
A.2 Details of text triplets
Here, we describe the details of the LLM-based and rule-based text triplet generation process. As shown in Figs. 4 and 5, which showcases examples of both LLM-based and rule-based triplets, both approaches produce natural and coherent text triplets. Note that none of the datasets used for generating text triplets overlap with the data used in the target CIR benchmarks, with the exception of the CASE dataset [22]. The source of the CASE dataset is VQA2.0 [15], which is constructed from the COCO dataset [27], potentially leading to overlap in cases involving COCO object composition [36].
[Detailed explanation on LLM-based triplets] As described in Sec. 3, besides Compodiff [16], we conduct experiments using various publicly available text triplets: IP2P [5], COVR [45], and CASE [22]. Although the primary objective of these approaches is to generate CIR triplets $(I_{r},T_{c},I_{t})$, they also produce text triplets. Below, we provide detailed descriptions of how text triplets are constructed in each approach (Note again that their final product is CIR triplets). There are two main ways to generate text triplets using LLMs: 1) generating both conditional text $T_{c}$ and target caption $T_{t}$ given reference caption $T_{r}$ using fine-tuned LLM for this task, such as IP2P, Compodiff; and 2) generating only conditional text $T_{t}$ given pairs $(T_{r},T_{t})$ from pre-existing captions by identifying with visually or text semantically similar such as CoVR [45], and CASE [22]. In addition to these existing datasets, we implement an efficient in-context learning-based generation process. Examples and summaries of each dataset can be found in Tab. 7 and Tab. 8.
IP2P employs GPT-3 for text triplets generation and fine-tunes it with a human-curated small set of 700 text triplets. Namely, given reference captions $T_{r}$ sampled from LAION-Aesthetics V2 6.5+ dataset [37], the corresponding conditional texts $T_{c}$ and corresponding target captions $T_{t}$ are manually written by humans. After fine-tuning on this small set of text triplets, the model generates 454k text triplets: reference captions $T_{r}$ from the LAION-Aesthetics V2 6.5+ dataset [37] are provided as input to the fine-tuned LLM, whose output predicts the corresponding conditioning text $T_{c}$ and the target caption $T_{t}$. Note that the LAION-Aesthetics dataset is not related to the original source datasets (FashionIQ, NLVR2, and MS-COCO) used in existing CIR benchmarks (FashionIQ, CIRR, and CIRCO), ensuring no overlap with the CIR benchmarks.
Compodiff enhances the scalability of the IP2P text triplet generation process by modifying the choice of LLM and expanding the fine-tuning dataset. As described in [[16], Section 4], the OPT-6.7B model is utilized and fine-tuned with LoRA on the above 454k text triplets of IP2P [5]. Then, similar to the IP2P approach, given reference captions from the LAION dataset [38], fine-tuned LLM generates the corresponding conditioning texts and target captions.
COVR starts by identifying similar caption pairs from the WebVid2M dataset [1], which contains 2.5 million video-caption pairs. These pairs serve as the reference captions ($T_{r}$) and target captions ($T_{t}$). Then, given these pairs ($T_{r},T_{t}$), LLM generates conditional captions that describe the differences between the paired captions. The LLaMA-7B model [43] is utilized for this purpose and is fine-tuned on an expanded version of the above 700 manually annotated triplets used in IP2P (adding 15 annotations for more diverse cases).
CASE uses VQA2.0 dataset [15], which consists of (image, question, answer) triplets. Given $(I,Q,A)$ triplets, complementary triplets $(I_{c},Q,A_{c})$ are manually selected based on visually similar image $I_{c}$ with three rules: 1) the premise assumed in question $Q$ holds for both $I$ and $I_{c}$, 2) $Q$ is logical for $I_{c}$, and 3) the answer $A_{c}$ for $I_{c}$ differs from $A$. Then, conditional text $T_{c}$ is generated by GPT-3, describing differences between image pairs $(I,I_{c})$ without fine-tuning, leveraging in-context learning with a few examples. Since the VQA2.0 dataset is derived from the COCO dataset, COCO captions that match VQA2.0 images are used to form text triplets.
As seen in Tab. 7, compared to other approaches, the quality of the relationships between $T_{r}$, $T_{c}$, and $T_{t}$ is not always satisfactory, which results in minimal performance gain as shown in Tab. 6. Namely, unlike other CIR datasets that first create high-quality text triplets before generating CIR triplets, CASE generates the conditioning text $T_{c}$ using the reference image $I_{r}$ and target image $I_{t}$. The provided reference text $T_{r}$ and target text $T_{t}$ are taken directly from the captions of reference image $I_{r}$ and target image $I_{t}$ in the VQA2.0 dataset. Therefore, due to the poor descriptiveness of these captions and their lack of consideration for the conditioning text $T_{c}$, while $T_{c}$ can effectively explain the visual differences between $I_{r}$ and $I_{t}$, it often fails to capture the differences between $T_{r}$ and $T_{t}$ adequately.
Efficient in-context learning refers to our efficient implementation which uses a recent and powerful LLM, LLaMA3-8B [14]. This approach performs in-context learning using reference captions $T_{r}$ from the CC3M dataset [39], guided by a custom-designed prompt with a few examples of textual modifications (e.g., replace, change, remove, …). Specifically, given a reference caption $T_{r}$, the prompt instructs the model to generate a target caption $T_{t}$, which is a complete sentence that slightly differs from the corresponding reference caption. Then, the prompt guides the model to generate conditioning text that explains the differences between $T_{r}$ and $T_{t}$, based on the above pre-defined textual modifications. Compared to Compodiff, which takes 3.8 hours to generate 1 million text triplets, this version requires only 1.5 hours. In Tab. 6, we verify that this more efficient version achieves competitive performance compared to the other fine-tuned LLM approaches.
Table 7: Examples of text triplet datasets.
Text triplets Reference text $T_{r}$ Conditioning text $T_{c}$ Target text $T_{t}$
Rule-based “another wall at my home” “bedroom is added in place of wall” “another bedroom at my home”
IP2P [5] “watercolor of your pet!” “make it a huge grizzly bear” “watercolor of a huge grizzly bear!”
Compodiff [16] “Chinese landscape watercolor painting” “make the landscape a cityscape” “chinese cityscape watercolor painting”
Efficient in-context learning “young business woman on a bench” “add a laptop” “young business woman on a bench with a laptop”
CoVR [45] “Two little boys are running” “Have them dance” “Two little boys are dancing”
CASE [22] “A scone with an orange slice on a plate” “This food is not acidic” “a close up of a muffin on a plate on a table”
Table 8: Summaries of text triplet dataset.
Dataset Use LLM Model Fine-tuning strategy # of text triplets
Rule-based 1.3M
IP2P [5] GPT-3 Fine-tuned on 700 human-written text triplets 450K
Compodiff [16] OPT-6.7B Fine-tuned on 450k IP2P text triplets 2.5M
Efficient in-context learning LLaMA3-8B In-context learning 1M
CoVR [45] LLaMA-7B Fine-tuned on 700 human-written text triplets 700K
CASE [22] GPT-3 In-context learning 350K
[Detailed explanation on rule-based triplets] To construct rule-based triplets, we mainly follow the process described in [[16], Section 4.1]. Firstly, given reference captions, important keywords like nouns are extracted with a part-of-speech (POS) tagger via the Spacy library. Then, the keyword is filtered by frequency filtering with hard thresholding to focus only on frequently occurring keywords. Specifically, we only use keywords that appear more than 100. After applying keyword frequency filtering, the remaining keyword list is used to create caption triplets $(T_{r},T_{c},T_{t})$. To generate text triplets, a keyword from the given $T_{r}$ is selected, and alternative keywords are extracted based on text similarity scores ranging from 0.5 to 0.7, using the SBERT all-MiniLM-L6-v2 [35]. The target caption $T_{t}$ is then constructed by substituting the original keyword with a similar alternative. The conditioning text $T_{c}$ is generated using randomly selected pre-defined templates, as detailed in Tab. 9. Here, most of the templates are similar to that of Compodiff [16]. We use captions from the CC3M dataset [39] as reference captions $T_{r}$. Note that CC3M is not related to the existing CIR benchmarks, which again ensures no overlap with the CIR benchmarks.
Since the quality of the generated triplets with the above procedure may not be optimal, we employ an additional filtering process. Compodiff [16] employs an additional filtering process that uses cosine similarities between generated images and texts, calculated by CLIP encoders. However, as we do not have images for captions, we filter the inappropriate texts using only textual information inspired by LinCIR [17]. Namely, we calculate the similarity between the CLIP text embedding of $T_{t}$ and the CLIP text embedding of “a photo of [\$currency-dollar\\$\$]” where [\$] is obtained by $T_{t}$ projected by $\phi$ from LinCIR (ViT-L/14). Following LinCIR noise ($\text{Unif}(0,1)\times\mathcal{N}(0,1)$) is injected before passing through $\phi$. After calculating the above similarity, texts whose similarities are less than the threshold (0.75) are removed. The same process is also applied to the reference caption $T_{r}$ and the intersection of filtering processes for $T_{t}$ and $T_{r}$ is used for the final dataset whose size becomes 1.3M. As described in Sec. B.6, we verify that this filtering process is effective. However, this does not imply that the effectiveness of rule-based text triplets is solely dependent on the use of a projection module in the filtering process; even without filtering, the enhancement from RTD remains significant.
Figure 4: Example of rule-based triplet datasets
Figure 5: Example of LLM-based triplet datasets
Table 9: The full 50 keyword converting templates
”replace \${source} with \${target} ”substitute \${target} for \${source}
”apply \${target} \${source} is removed and \${target} takes its place”
”convert \${source} to \${target} ”modify \${source} to become \${target}
”replace \${source} with \${target} ”customize \${source} to become \${target}
”update \${source} to \${target} ”change \${source} to match \${target}
”substitute \${target} for \${source} \${target} is introduced after \${source} is removed”
”alter \${source} to match \${target} \${target} is added in place of \${source}
”upgrade \${source} to \${target} \${target} is introduced as the new option after”
”amend \${source} to fit \${target} \${source} is removed and \${target} is added”
”opt for \${target} \${source} is removed and \${target} is introduced”
\${source} is removed” \${target} is added as a replacement for \${source}
”add \${target} \${target} is the new option available”
”if it is \${target} \${target} is added after \${source} is removed”
\${target} is the updated option” \${target} is introduced after \${source} is retired”
\${target} is the updated choice” ”tweak \${source} to become \${target}
\${source} is replaced with \${target} ”has no \${source}
”change \${source} to \${target} ”alter \${source} to \${target}
”swap \${source} for \${target} ”redesign \${source} as \${target}
”turn \${source} into \${target} ”adapt \${source} to fit \${target}
”choose \${target} instead of \${source} \${target} is the new choice”
\${target} is the new selection” ”exchange \${source} with \${target}
”transform \${source} into \${target} ”show no \${source}
”no \${source} ”remove \${source}
”delete \${source} ”not a \${source}
”with no \${source} ”without \${source}
A.3 Details on integration with FTI4CIR [26] and Context-I2W [42].
FTI4CIR [26] enhances the fine-grained capability of the projection module $\phi$ by separately handling subjects and attributes using distinct projection modules, $\phi_{s}$ and $\phi_{a}$. To achieve this, they leverage BLIP-generated captions that explicitly separate subject and attribute information in the text domain, formatted as “[primary subject(s)] + [detailed description].” Unlike the subject module $\phi_{s}$, which focuses on global subject information, $\phi_{a}$ processes localized attribute details. FTI4CIR extracts attribute features by adaptively aggregating patch features via an additional transformer. However, this approach requires an extra forward pass of the transformer with full sequences of visual embeddings for $\phi_{a}$, rather than a single pooled embedding, making it incompatible with our refined concatenation scheme (RC). Therefore, we do not use $\phi_{a}$ during the training of RTD (of course, it is used in inference). Instead, we focus on training $\phi_{s}$ while incorporating attribute information in the text domain. Specifically, we generate text triplets using BLIP-generated captions to separately capture subject and attribute information, through “Efficient In-Context Learning” (see Sec. A.2). Then, we provide subject information to $\phi_{s}$ while directly using attribute information in text format. For example, given text triplets: $T_{r}:$ “a room with a chair and a table”, $T_{c}:$ “replace the chair with a sofa”, $T_{t}:$ “a room with a sofa and a table”. our refined concatenation scheme (RC) ensures that only textual subject information is input to $\phi_{s}$. This extraction is feasible because BLIP-generated captions (from FTI4CIR) already provide subject and attribute separation.
Context-I2W [42] introduces a context-dependent projection module that selects relevant visual information. To achieve this, Context-I2W employs a more complex projection mechanism that utilizes context captions such as “A [REPLACE] passes the ball with his teammate during a training session”, which removes the first subject term. Then, these context captions and full visual sequence embeddings from the visual encoder are fused in the Context-I2W projection module. Since it requires full sequences of visual embeddings rather than a single pooled embedding, it can be incompatible with our refined concatenation scheme (RC). In such cases, our method can still be applied simply by replacing the text encoder. Namely, instead of training RTD with this projection module, we simply replace the text encoder with the one updated by RTD using the projection module $\phi$ from Pic2Word [36]. Namely, during inference, we use the projection module from Context-I2W alongside the RTD-updated text encoder (which was trained with the projection module from Pic2Word). The strong performance of this variant suggests that RTD does not need to be trained with each specific projection module, again highlighting its strong generalizability.
Appendix B Additional experimental results
B.1 Results on GeneCIS [44] and COCO object composition
We observe that integrating our approach with projection-based CIR methods results in consistent yet marginal performance improvements on GeneCIS, as shown in Tab. 10. The relatively smaller performance difference compared to other datasets can be attributed to the discrepancy between the format of the conditioning text of GeneCIS and the projection-based CIR methods training methodology. Namely, GeneCIS only uses the fixed four text instructions “change attribute”, “focus attribute”, “change object” and “focus object”, which is different from the usual text instruction we expected (e.g., “change the dog to a cat”).
In the experiment on COCO object composition, we observe a significant performance improvement, similar to the results obtained on other datasets in Tab. 11. This finding reaffirms that our approach, when combined with ZS-CIR methods, consistently achieves strong performance, demonstrating its generalizability.
Table 10: GeneCIS results
Average
R@1 R@2 R@3
ViT-B Pic2Word $11.13$ $21.08$ $31.05$
SEARLE $12.19$ $22.56$ $32.03$
+RTD $12.82$ ($\mathbf{+0.63}$) $22.97$ ($\mathbf{+0.41}$) $32.44$ ($\mathbf{+0.41}$)
LinCIR $12.23$ $21.29$ $30.80$
+RTD $12.83$ ($\mathbf{+0.60}$) $22.83$ ($\mathbf{+1.54}$) $32.22$ ($\mathbf{+1.42}$)
ViT-L Pic2Word $11.18$ $21.45$ $30.55$
+RTD $11.92$ ($\mathbf{+0.74}$) $22.32$ ($\mathbf{+0.87}$) $31.33$ ($\mathbf{+0.78}$)
SEARLE $12.30$ $22.08$ $31.29$
+RTD $12.40$ ($\mathbf{+0.10}$) $22.82$ ($\mathbf{+0.74}$) $32.37$ ($\mathbf{+1.08}$)
LinCIR $12.45$ $22.66$ $32.06$
+RTD $13.18$ ($\mathbf{+0.73}$) $23.12$ ($\mathbf{+0.46}$) $32.77$ ($\mathbf{+0.71}$)
Table 11: COCO object composition results
COCO
R@1 R@5 R@10
ViT-B Pic2Word $6.88$ $13.6$ $17.52$
+RTD $7.62$ ($\mathbf{+0.74}$) $20.23$ ($\mathbf{+6.63}$) $28.69$ ($\mathbf{+11.17}$)
SEARLE $9.52$ $21.45$ $29.38$
+RTD $11.01$ ($\mathbf{+1.49}$) $24.34$ ($\mathbf{+2.89}$) $32.84$ ($\mathbf{+3.46}$)
LinCIR $7.15$ $18.38$ $27.3$
+RTD $9.59$ ($\mathbf{+2.44}$) $21.66$ ($\mathbf{+3.28}$) $30.66$ ($\mathbf{+3.36}$)
ViT-L Pic2Word $10.26$ $23.67$ $32.53$
+RTD $10.26$ ($\mathbf{+0.00}$) $24.66$ ($\mathbf{+0.99}$) $33.56$ ($\mathbf{+1.03}$)
SEARLE $12.07$ $26.13$ $35.17$
+RTD $14.38$ ($\mathbf{+2.31}$) $29.74$ ($\mathbf{+3.61}$) $38.09$ ($\mathbf{+2.92}$)
LinCIR $11.37$ $24.53$ $33.85$
+RTD $14.6$ ($\mathbf{+3.23}$) $29.84$ ($\mathbf{+5.31}$) $38.87$ ($\mathbf{+5.02}$)
B.2 Results on different backbones (ViT-B/32, ViT-L/14)
Tab. 12 summarizes the evaluation results on the FashionIQ dataset. In the table, we observe that the incorporation of our approach with projection-based CIR methods significantly improves the performance across all three existing projection-based CIR methods (SEARLE, Pic2Word, and LinCIR) and all backbones (ViT-B/32 and ViT-L/14). For example, regardless of the choice of projection-based CIR methods and backbones, the minimum performance gain for average R@10 and R@50 scores is greater than 2 and 3.5 points, respectively. Tab. 13 shows a similar trend on the CIRR and CIRCO datasets. Notably, in some metrics on the CIRR and CIRCO datasets, the performance improvements achieved through our method (ViT-B/32) even exceed those obtained by employing a larger backbone (ViT-L/14), which demonstrates the effect of our method. Specifically, in the CIRR R@1 score, SEARLE + RTD (26.29) and LinCIR + RTD (24.82) using ViT-B/32 surpasses the original results of SEARLE (24.89) and LinCIR (23.76) using ViT-L/14.